I know you have to square at least twice but it get really confusing. Please help:
√x - √x-5 =1


Answer by rothauserc(4718)

solve sqrt(x) -sqrt(x-5) = 1
add sqrt(x-5) to both sides of =
sqrt(x) = (sqrt(x-5)+1)
now square both sides of =
x = x-5 +2*sqrt(x-5) +1
subtract x from both sides of =
0 = -4 +2*sqrt(x-5)
add -4 to both sides of =
4 = 2*sqrt(x-5)
divide both sides of = by 2
2 = sqrt(x-5)
square both sides of =
4 = x-5
===================================
x = 9 is the answer!
==================================
check in original equation
sqrt(9) - sqrt(9-5) = 1
3 - 2 = 1
our answer checks!